hwo wold you solove 7 exponent 4 over 7 exponent 8
Answer by stanbon(75887)

solve 7 exponent 4 over 7 exponent 8
----
The exponent tells you the number of
times the base number is used as a factor.
---
Your Problem:
[7*7*7*7[/[7*7*7*7*7*7*7*7]
---
Cancel as many common factors as you can and you will get:
---
= 1/[7*7*7*7]
---
= 1/7^4
----------------
The exponent law does the problem as follows:
7^4/7^8 = 1/[7^(8-4)] = 1/[7^4]
=======================================
Cheers,
Stan H.
============